OpenStudy (anonymous):

I just need help setting up the equation will give medal and become a fan!!!!!!! The art museum had a total of 224 visitors on Tuesday. Visitors older than 18 paid $12 for admission. Visitors 18 years or younger paid $8 for admission. The museum collected a total of $2,520. How many visitors were 18 or younger? 42 56 182 194

OpenStudy (dbzfan836):

hmm, well, you could guess and check

OpenStudy (dbzfan836):

but that could take a long period of time

OpenStudy (dbzfan836):

Wait, i have an answer

OpenStudy (dbzfan836):

A 42

OpenStudy (dbzfan836):

hmm, well that didnt take long at all XD much shorter time used than expected

OpenStudy (anonymous):

let x be visitors under 18 years age=x,then \[8x+(224-x)*12=2520\] 8x+2688-12x=2520 -4x=2520-2688 x=?

OpenStudy (dbzfan836):

X=42, then, and that there assures it

OpenStudy (dbzfan836):

but you are missing a step, @surjithayer

OpenStudy (dbzfan836):

wait, ALL the fours in the semifinal line are negative, I made a mishap

OpenStudy (dbzfan836):

8x+2688-12x=2520 -4x=2520-2688 -4x/-4=-166/-4 x=?

OpenStudy (anonymous):

So what would be the process to solve if@surjithayer is missing a step?!

OpenStudy (anonymous):

i have short listed in line two.i.e, visitors over 18=224-x in first line it should be visitors of 18 and under 18=x

OpenStudy (anonymous):

-4x=-168 \[x=\frac{ -168 }{ -4 }=42\]

OpenStudy (anonymous):

Ok thank you guys so much !

OpenStudy (anonymous):

yw